# Andreas Schlaginhaufen SYCAMORE, EPFL

### v1

@@ 3.3 Theoretical insights
propositionbregmandivergence Under Assumptions 2 and 2, we have $\ell(r^{\prime},\mu)=D_{\bar{h}}(\mu,\mathsf{RL}(r^{\prime}))$ for any $\mu\in\mathcal{M}$. Proposition 10 above demonstrates that the suboptimality of an occupancy measure $\mu$ for the reward $r^{\prime}$ coincides with the Bregman divergence between $\mu$ and the optimal occupancy measure under $r^{\prime}$. This generalizes (Mei et al., 2020, Lemma 26) from entropy regularization to any steeply regularized MDP. The proof is presented in Appendix D.6.
@@
$\frac{\sigma_{\mathcal{R}}}{2}\left\lVert[r]_{\mathcal{U}}-[r^{\prime}]_{\mathcal{U}}\right\rVert_{2}^{2}\leq\ell(r^{\prime},\mathsf{RL}(r))=D_{\bar{h}}\left({\mathsf{RL}(r),\mathsf{RL}(r^{\prime})}\right)\leq\frac{1}{2\eta}\left\lVert[r]_{\mathcal{U}}-[r^{\prime}]_{\mathcal{U}}\right\rVert^{2}_{2},$ (13)
@@
Example 3.3 (continues=ex:1).
Recall the definition $\mathcal{U}_{P^{k}}=\operatorname{im}(E-\gamma P^{k})$. Given that in Example 3.1 we have $\ell_{P^{0}}(\hat{r},\mu^{\textsf{E}}_{P^{0}})=0$ and $\ell_{P^{1}}(\hat{r},\mu^{\textsf{E}}_{P^{1}})=\mathcal{O}(\beta)$, Lemma 3.3 ensures that $\hat{r}$ and $r^{\textsf{E}}$ coincide in $\mathbb{R}^{\mathcal{S}\times\mathcal{A}}/\mathcal{U}_{P^{0}}$, and for small $\beta$, they are close in $\mathbb{R}^{\mathcal{S}\times\mathcal{A}}/\mathcal{U}_{P^{1}}$. However, as illustrated in Figure1(a) this doesn’t ensure that $\hat{r}$ and $r^{\textsf{E}}$ are close in $\mathbb{R}^{\mathcal{S}\times\mathcal{A}}/\bm{1}$ and $\mathbb{R}^{\mathcal{S}\times\mathcal{A}}/\mathcal{U}_{P}$. In particular, it can be computed that $\left\lVert[\hat{r}]_{\mathcal{U}_{P}}-[r^{\textsf{E}}]_{\mathcal{U}_{P}}\right\rVert_{2}\approx 1.51$, which by Lemma 3.3 explains the poor transferability to $P$.
(a) Rewards in Example 3.1.                            (b) Proof sketch Theorem 3.4.
Figure 1: For a small $\beta$, (a) illustrates the equivalence classes $[\hat{r}]_{\mathcal{U}}$ and $[r^{\textsf{E}}]_{\mathcal{U}}$, corresponding to the transition laws $P^{0},P^{1},P$ from Example 3.1, in $\mathbb{R}^{\mathcal{S}\times\mathcal{A}}/\bm{1}$. The blue lines correspond to $P^{0}$, the red lines to $P^{1}$, and the gray lines to $P$. Furthermore, the shaded areas illustrate the approximation error around $[r^{\textsf{E}}]_{\mathcal{U}_{P^{k}}}$, as guaranteed by Lemma 3.3. (b) illustrates the uncertainty set for the recovered reward when learning from two experts, as discussed in the proof sketch of Theorem 3.4.
@@ 3.4 Sufficient conditions for transferability
Some observations are in order. First, the above theorem shows that the larger the second principal angle between the two experts’ transition laws, the better the recovered reward transfers to a new environment. Second, observe that $\theta_{2}({P^{0}},{P^{1}})>0$ is equivalent to the rank condition (9), as the second principal angle between two subspaces is non-zero if and only if their intersection is at most one-dimensional. Therefore, for exact transferability, Theorem 3.4 requires the rank condition (9) to be satisfied and $\hat{\varepsilon}=0$, recovering the results by Cao et al. (2021); Rolland et al. (2022); Schlaginhaufen and Kamgarpour (2023). However, in contrast to past results, Theorem 3.4 applies to more realistic scenarios, where $\hat{\varepsilon}$ is merely small, not zero. Finally, we note that Theorem 3.4 can be trivially generalized to $K\geq 2$ experts by replacing $\theta_{2}({P^{0}},{P^{1}})$ with the maximum of $\theta_{2}({P^{k}},{P^{l}})$ over $10\leq k\leq l\leq K-$ 1.
@@
When learning a reward $\hat{r}$ from a single expert ($K=1$), Schlaginhaufen and Kamgarpour (2023) show that, without reducing the dimension of the reward class, $\hat{r}$ cannot be exactly transferable to any neighborhood of the expert’s transition law $P_{0}$. However, Theorem 3.4 below shows that by allowing for an $\varepsilon$ of error, we can guarantee transferability to a neighborhood of $P_{0}$.{restatable}theoremlocaltransferability Let $K=1$, $D\vcentcolon=\max_{r,r^{\prime}\in\mathcal{R}}\left\lVert r-r^{\prime}\right\rVert_{2}$, and suppose that Assumptions 2,2, and 3.3 hold. If $\ell_{P^{0}}(\hat{r},\mu^{\textsf{E}})\leq\hat{\varepsilon}$, then $\hat{r}$ is $\varepsilon_{P}$-transferable to $P\in\Delta_{\mathcal{S}}^{\mathcal{S}\times\mathcal{A}}$ with
@@
To provide more insights about Theorems 3.4 and 3.4, we provide explicit values for the primal and dual strong convexity constants, $\eta$ and $\sigma_{\mathcal{R}}$, respectively. To this end, we focus on the Shannon entropy regularization $h(p)=-\tau\mathcal{H}(p)$ and the Tsallis-1/2 entropy regularization $h(p)=-\tau\mathcal{H}_{1/2}(p)$ as defined in Appendix C. While the Shannon entropy regularization is commonly used in IRL (Ziebart, 2010; Ho and Ermon, 2016), the Tsallis-1/2 entropy is more often adopted in the multi-armed bandit literature Zimmert and Seldin (2021). Both regularizations satisfy Assumption 2 as well as Assumption 3.3 with the constants detailed in Proposition D.5 in the appendix. In general, the Tsallis entropy leads to a slightly smaller strong convexity constant $\eta$, but avoids an exponential dependence on the effective horizon $H_{\gamma}=1/(1-\gamma)$ in $\sigma_{\mathcal{R}}$. Below, we summarize the implications of Proposition D.5 for $\varepsilon$-transferability of a reward $\hat{r}$ recovered from two experts. {restatable}corollaryexplicit_transferability Suppose the conditions in Theorem 3.4 hold. Furthermore, let $H_{\gamma}\vcentcolon=1/(1-\gamma)$, $R\vcentcolon=\max_{r\in\mathcal{R}}\left\lVert r\right\rVert_{\infty}$, $D=\max_{r,r^{\prime}\in\mathcal{R}}\left\lVert r-r^{\prime}\right\rVert_{2}$, and $\tau<2D$. Then, for the Shannon entropy $\hat{r}$ is $\varepsilon$-transferable to $\mathcal{P}=\Delta_{\mathcal{S}}^{\mathcal{S}\times\mathcal{A}}$ with
$\varepsilon=\dfrac{H_{\gamma}^{2}D|\mathcal{S}||\mathcal{A}|^{2+H_{\gamma}}\exp\left({\frac{2RH_{\gamma}}{\tau}}\right)}{\nu_{\text{min}}^{2}\tau\left({1-\frac{\tau}{2D}}\right)\sin\left({\theta_{2}({P^{0}},{P^{1}})/2}\right)^{2}}\hat{\varepsilon},$ (14)
@@
$\varepsilon=\dfrac{2\sqrt{2}H_{\gamma}^{5}D|\mathcal{S}||\mathcal{A}|^{2}\left({2R/\tau+3\sqrt{|\mathcal{A}|}}\right)^{3}}{\nu_{\text{min}}^{2}\tau\left({1-\frac{\tau}{2D}}\right)\sin\left({\theta_{2}({P^{0}},{P^{1}})/2}\right)^{2}}\hat{\varepsilon}.$ (15)
@@
Our results in this section, especially Proposition 10 and Lemma 3.3, are critically relying on the steepness of the regularization (Assumption 2), which is essential to ensure that the expert’s reward can be identified up to the equivalence class of potential shaping transformations. Although we can still upper bound the suboptimality $\ell(r^{\textsf{E}},\mathsf{RL}(\hat{r}))$ in terms of the distance between $\hat{r}$ to $r^{\textsf{E}}$ in $\mathbb{R}^{\mathcal{S}\times\mathcal{A}}/\mathcal{U}$ without this assumption (see Proposition D.6), we no longer have a lower bound as in Lemma 3.3, which is essential for establishing closeness of $\hat{r}$ and $r^{\textsf{E}}$ in $\mathbb{R}^{\mathcal{S}\times\mathcal{A}}/\mathcal{U}$. Hence, we expect it to be difficult to obtain guarantees similar to those in Theorem 3.4 and 3.4 for the unregularized setting, without either reducing the dimension of the reward class (Amin et al., 2017) or making specific assumptions about the feasible reward sets (Metelli et al., 2021, Assumption 4.1).
@@ 4 Algorithm
To provide end-to-end guarantees for recovering transferable rewards from a finite set of expert demonstrations, we analyze the convergence and sample complexity (in terms of expert demonstrations) of an algorithm for recovering a reward for which all $K$ experts are approximately optimal. To this end, we focus on the reward class =$\mathcal{R}=\left\{{r\in\mathbb{R}^{\mathcal{S}\times\mathcal{A}}:\left\lVert r\right\rVert_{1}\leq 1}\right\}$. Furthermore, we assume oracle access to a $(\varepsilon,\delta)$-PAC algorithm for the forward problem (O-RL). That is, a polynomial-time algorithm, $\mathsf{A}^{\varepsilon,\delta}$, that outputs a policy $\pi=\mathsf{A}^{\varepsilon,\delta}(r)$ such that with probability at least $1-\delta$ it holds that $\ell(r,\mu^{\pi})\leq\varepsilon$ (see e.g. (Lan, 2023) for a specific example). The key idea of our meta-algorithm is to learn a reward minimizing the sum of the suboptimalities of the $K$ experts $\mu^{\textsf{E}}_{P^{0}},\ldots,\mu^{\textsf{E}}_{P^{K-1}}$. This leads us to the following multi-expert IRL problem
$\min_{r\in\mathcal{R}}\sum_{k=0}^{K-1}\ell_{P^{k}}(r,\hat{\mu}_{\mathcal{D}^{\textsf{E}}_{k}}),$ (O-IRL)
@@
theoremirlconvergence Suppose that $N^{\textsf{E}}=\Omega\big{(}K\log(|\mathcal{S}||\mathcal{A}|/\hat{\delta})/\hat{\varepsilon}^{2}\big{)}$ and $H^{\textsf{E}}=\Omega\big{(}\log(K/\hat{\varepsilon})/\log(1/\gamma)\big{)}$. Running Algorithm 1 for $T=\Omega\big{(}K^{2}/\hat{\varepsilon}^{2}\big{)}$ iterations with step-size $\alpha=1/(K\sqrt{T})$, where $\delta_{\text{opt}}=\mathcal{O}\big{(}\hat{\delta}\hat{\varepsilon}^{2}/K^{3}\big{)}$ , $\varepsilon_{\text{opt}}=\mathcal{O}(\hat{\varepsilon}/K)$, $N=\Omega\big{(}K\log(K|\mathcal{S}||\mathcal{A}|/(\hat{\delta}\hat{\varepsilon}))/\hat{\varepsilon}^{2}\big{)}$, and $H=H^{\textsf{E}}$, it holds with probability at least $1-\hat{\delta}$ that $\ell_{P^{k}}(\hat{r},\mu^{\textsf{E}}_{P^{k}})\leq\hat{\varepsilon},\;\text{ for }k=0,\ldots,K-1.$ The above result generalizes (Syed and Schapire, 2007, Theorem 2) by considering multiple experts and by proving convergence in terms of the expert suboptimality. We refer to Appendix H for the proof and the precise constants. Theorem 1 shows that with $\Omega(K/\hat{\varepsilon}^{2})$ demonstrations of each expert, we recover in $\Omega(K^{2}/\hat{\varepsilon}^{2})$ steps of Algorithm 1 a reward $\hat{r}$ for which all experts are $\hat{\varepsilon}$-optimal. Together with Theorem 3.4 and 3.4, this provides a bound on the sample and time complexity of recovering in $\varepsilon$-transferable rewards in regularized IRL.

### v2

@@ 3.3 Theoretical insights
propositionbregmandivergence Under Assumptions 2 and 2, we have $\mathsf{SubOpt}(r^{\prime},\mu)=D_{\bar{h}}(\mu,\mathsf{RL}(r^{\prime}))$ for any $\mu\in\mathcal{M}$. Proposition 10 above demonstrates that the suboptimality of an occupancy measure $\mu$ for the reward $r^{\prime}$ coincides with the Bregman divergence between $\mu$ and the optimal occupancy measure under $r^{\prime}$. This generalizes (Mei et al., 2020, Lemma 26) from entropy regularization to any steeply regularized MDP. The proof is presented in Appendix D.6.
@@
$\frac{\sigma_{\mathcal{R}}}{2}\left\lVert[r]_{\mathcal{U}}-[r^{\prime}]_{\mathcal{U}}\right\rVert_{2}^{2}\leq\mathsf{SubOpt}(r^{\prime},\mathsf{RL}(r))=D_{\bar{h}}\left({\mathsf{RL}(r),\mathsf{RL}(r^{\prime})}\right)\leq\frac{1}{2\eta}\left\lVert[r]_{\mathcal{U}}-[r^{\prime}]_{\mathcal{U}}\right\rVert^{2}_{2},$ (13)
@@
Example 3.4 (continues=ex1).
Recall the definition $\mathcal{U}_{P^{k}}=\operatorname{im}(E-\gamma P^{k})$. Given that in Example 3.2 we have $\mathsf{SubOpt}_{P^{0}}(\hat{r},\mu^{\textsf{E}}_{P^{0}})=0$ and $\mathsf{SubOpt}_{P^{1}}(\hat{r},\mu^{\textsf{E}}_{P^{1}})=\mathcal{O}(\beta)$, Lemma 3.3 ensures that $\hat{r}$ and $r^{\textsf{E}}$ coincide in $\mathbb{R}^{\mathcal{S}\times\mathcal{A}}/\mathcal{U}_{P^{0}}$, and for small $\beta$, they are close in $\mathbb{R}^{\mathcal{S}\times\mathcal{A}}/\mathcal{U}_{P^{1}}$. However, as illustrated in Figure1(a) this doesn’t ensure that $\hat{r}$ and $r^{\textsf{E}}$ are close in $\mathbb{R}^{\mathcal{S}\times\mathcal{A}}/\bm{1}$ and $\mathbb{R}^{\mathcal{S}\times\mathcal{A}}/\mathcal{U}_{P}$. In particular, it can be computed that $\left\lVert[\hat{r}]_{\mathcal{U}_{P}}-[r^{\textsf{E}}]_{\mathcal{U}_{P}}\right\rVert_{2}\approx 1.51$, which by Lemma 3.3 explains the poor transferability to $P$.
@@ 3.4 Sufficient conditions for transferability
(a) Rewards in Example 3.2.                            (b) Proof sketch Theorem 3.4.
Figure 1: (a) illustrates the equivalence classes $[\hat{r}]_{\mathcal{U}}$ and $[r^{\textsf{E}}]_{\mathcal{U}}$, corresponding to the transition laws $P^{0},P^{1},P$ from Example 3.2, for a small $\beta$, in $\mathbb{R}^{\mathcal{S}\times\mathcal{A}}/\bm{1}$. The blue lines correspond to $P^{0}$, the red lines to $P^{1}$, and the gray lines to $P$. Furthermore, the shaded areas illustrate the approximation error around $[r^{\textsf{E}}]_{\mathcal{U}_{P^{k}}}$, as guaranteed by Lemma 3.3. (b) illustrates the uncertainty set for the recovered reward when learning from two experts, as discussed in the proof sketch of Theorem 3.4.
Some observations are in order. First, the above theorem shows that the larger the second principal angle between the two experts’ transition laws, the better the recovered reward transfers to a new environment. Second, observe that $\theta_{2}({P^{0}},{P^{1}})>0$ is equivalent to the rank condition (9), as the second principal angle between two subspaces is non-zero if and only if their intersection is at most one-dimensional. Therefore, for exact transferability, Theorem 3.4 requires the rank condition (9) to be satisfied and $\hat{\varepsilon}=0$, recovering the results by Cao et al. (2021); Rolland et al. (2022); Schlaginhaufen and Kamgarpour (2023). But in contrast to past results, Theorem 3.4 applies to more realistic scenarios, where $\hat{\varepsilon}$ is merely small, not zero. Finally, we note that Theorem 3.4 can be trivially generalized to $K\geq 2$ experts by replacing $\theta_{2}({P^{0}},{P^{1}})$ with the maximum of $\theta_{2}({P^{k}},{P^{l}})$ over $10\leq k\leq l\leq K-$ 1. However, such bounds may be loose for $K>2$, potentially leaving considerable room for improvement in this setting.
@@
When learning a reward $\hat{r}$ from a single expert ($K=1$), Schlaginhaufen and Kamgarpour (2023) show that, without reducing the dimension of the reward class, $\hat{r}$ cannot be exactly transferable to any neighborhood of the expert’s transition law $P_{0}$. However, Theorem 3.4 below shows that by allowing for an $\varepsilon$ of error, we can guarantee transferability to a neighborhood of $P_{0}$.{restatable}theoremlocaltransferability Let $K=1$, $D\vcentcolon=\max_{r,r^{\prime}\in\mathcal{R}}\left\lVert r-r^{\prime}\right\rVert_{2}$, and suppose that Assumptions 2,2, and 3.3 hold. If $\mathsf{SubOpt}_{P^{0}}(\hat{r},\mu^{\textsf{E}})\leq\hat{\varepsilon}$, then $\hat{r}$ is $\varepsilon_{P}$-transferable to $P\in\Delta_{\mathcal{S}}^{\mathcal{S}\times\mathcal{A}}$ with
@@
Remark 3.5.
As discussed in Appendix H, we can compute the principal angles using a singular value decomposition. Moreover, given estimates $\hat{P}^{0},\hat{P}^{1}$ of the transition laws $P^{0},P^{1}$, the error in the estimate of $\sin\theta_{i}(P^{0},P^{1})$ scales with $\mathcal{O}(\max\{||P^{0}-\hat{P}^{0}||,||P^{1}-\hat{P}^{1}||\})$.
@@
To provide more insights about Theorems 3.4 and 3.4, we provide explicit values for the primal and dual strong convexity constants, $\eta$ and $\sigma_{\mathcal{R}}$, respectively. To this end, we focus on the Shannon entropy regularization $h(p)=-\tau\mathcal{H}(p)$ and the Tsallis-1/2 entropy regularization $h(p)=-\tau\mathcal{H}_{1/2}(p)$ as defined in Appendix C. While the Shannon entropy regularization is commonly used in IRL (Ziebart, 2010; Ho and Ermon, 2016), the Tsallis-1/2 entropy is more often adopted in the multi-armed bandit literature Zimmert and Seldin (2021). Both regularizations satisfy Assumption 2 as well as Assumption 3.3 with the constants detailed in Proposition D.5 in the appendix. In general, the Tsallis entropy leads to a slightly smaller strong convexity constant $\eta$, but avoids an exponential dependence on the effective horizon $H_{\gamma}=1/(1-\gamma)$ in $\sigma_{\mathcal{R}}$. Below, we summarize the implications of Proposition D.5 for $\varepsilon$-transferability of a reward $\hat{r}$ recovered from two experts. {restatable}corollaryexplicit_transferability Suppose the conditions in Theorem 3.4 hold. Furthermore, let $H_{\gamma}\vcentcolon=1/(1-\gamma)$, $R\vcentcolon=\max_{r\in\mathcal{R}}\left\lVert r\right\rVert_{\infty}$, $D=\max_{r,r^{\prime}\in\mathcal{R}}\left\lVert r-r^{\prime}\right\rVert_{2}$, and $\tau<D$. Then, for the Shannon entropy $\hat{r}$ is $\varepsilon$-transferable to $\mathcal{P}=\Delta_{\mathcal{S}}^{\mathcal{S}\times\mathcal{A}}$ with
$\varepsilon=\dfrac{2H_{\gamma}^{2}D|\mathcal{S}||\mathcal{A}|^{2+H_{\gamma}}\exp\left({\frac{2RH_{\gamma}}{\tau}}\right)}{\nu_{\text{min}}^{2}\tau\sin\left({\theta_{2}({P^{0}},{P^{1}})/2}\right)^{2}}\hat{\varepsilon},$ (14)
@@
$\varepsilon=\dfrac{4\sqrt{2}H_{\gamma}^{5}D|\mathcal{S}||\mathcal{A}|^{2}\left({2R/\tau+3\sqrt{|\mathcal{A}|}}\right)^{3}}{\nu_{\text{min}}^{2}\tau\sin\left({\theta_{2}({P^{0}},{P^{1}})/2}\right)^{2}}\hat{\varepsilon}.$ (15)
@@
Our results in this section, especially Proposition 10 and Lemma 3.3, are critically relying on the steepness of the regularization (Assumption 2), which is essential to ensure that the expert’s reward can be identified up to the equivalence class of potential shaping transformations. Although we can still upper bound the suboptimality $\mathsf{SubOpt}(r^{\textsf{E}},\mathsf{RL}(\hat{r}))$ in terms of the distance between $\hat{r}$ to $r^{\textsf{E}}$ in $\mathbb{R}^{\mathcal{S}\times\mathcal{A}}/\mathcal{U}$ without this assumption (see Proposition D.7), we no longer have a lower bound as in Lemma 3.3, which is essential for establishing closeness of $\hat{r}$ and $r^{\textsf{E}}$ in $\mathbb{R}^{\mathcal{S}\times\mathcal{A}}/\mathcal{U}$. Hence, we expect it to be difficult to obtain guarantees similar to those in Theorem 3.4 and 3.4 for the unregularized setting, without either reducing the dimension of the reward class (Amin et al., 2017) or making specific assumptions about the feasible reward sets (Metelli et al., 2021, Assumption 4.1).
@@ 4 Algorithm
To provide end-to-end guarantees for recovering transferable rewards from a finite set of expert demonstrations, we analyze the convergence and sample complexity (in terms of expert demonstrations) of an algorithm for recovering a reward for which all $K$ experts are approximately optimal. To this end, we focus on the reward class =$\mathcal{R}=\left\{{r\in\mathbb{R}^{\mathcal{S}\times\mathcal{A}}:\left\lVert r\right\rVert_{1}\leq 1}\right\}$. Furthermore, we assume oracle access to a $(\varepsilon,\delta)$-PAC algorithm for the forward problem (O-RL). That is, a polynomial-time algorithm, $\mathsf{A}^{\varepsilon,\delta}$, that outputs a policy $\pi=\mathsf{A}^{\varepsilon,\delta}(r)$ such that with probability at least $1-\delta$ it holds that $\mathsf{SubOpt}(r,\mu^{\pi})\leq\varepsilon$ (see e.g. (Lan, 2023) for a specific example). The key idea of our meta-algorithm is to learn a reward minimizing the sum of the suboptimalities of the $K$ experts $\mu^{\textsf{E}}_{P^{0}},\ldots,\mu^{\textsf{E}}_{P^{K-1}}$. This leads us to the following multi-expert IRL problem
$\min_{r\in\mathcal{R}}\sum_{k=0}^{K-1}\mathsf{SubOpt}_{P^{k}}(r,\hat{\mu}_{\mathcal{D}^{\textsf{E}}_{k}}),$ (O-IRL)
@@
theoremirlconvergence Suppose that $N^{\textsf{E}}=\Omega\big{(}K\log(|\mathcal{S}||\mathcal{A}|/\hat{\delta})/\hat{\varepsilon}^{2}\big{)}$ and $H^{\textsf{E}}=\Omega\big{(}\log(K/\hat{\varepsilon})/\log(1/\gamma)\big{)}$. Running Algorithm 1 for $T=\Omega\big{(}K^{2}/\hat{\varepsilon}^{2}\big{)}$ iterations with step-size $\alpha=1/(K\sqrt{T})$, where $\delta_{\text{opt}}=\mathcal{O}\big{(}\hat{\delta}\hat{\varepsilon}^{2}/K^{3}\big{)}$ , $\varepsilon_{\text{opt}}=\mathcal{O}(\hat{\varepsilon}/K)$, $N=\Omega\big{(}K\log(K|\mathcal{S}||\mathcal{A}|/(\hat{\delta}\hat{\varepsilon}))/\hat{\varepsilon}^{2}\big{)}$, and $H=H^{\textsf{E}}$, it holds with probability at least $1-\hat{\delta}$ that $\mathsf{SubOpt}_{P^{k}}(\hat{r},\mu^{\textsf{E}}_{P^{k}})\leq\hat{\varepsilon},\;\text{ for }k=0,\ldots,K-1.$ The above result generalizes (Syed and Schapire, 2007, Theorem 2) by considering multiple experts and by proving convergence in terms of the expert suboptimality. We refer to Appendix I for the proof and the precise constants. Theorem 1 shows that with $\Omega(K/\hat{\varepsilon}^{2})$ demonstrations of each expert, we recover in $\Omega(K^{2}/\hat{\varepsilon}^{2})$ steps of Algorithm 1 a reward $\hat{r}$ for which all experts are $\hat{\varepsilon}$-optimal. Together with Theorem 3.4 and 3.4, this provides a bound on the sample and time complexity of recovering in $\varepsilon$-transferable rewards in regularized IRL.
(a)                         (b)                         (c)                         (d)
Figure 2: (a) shows the second principal angle between the experts, for varying wind strength $\beta$. (b) shows the distance between $\hat{r}$ and $r^{\textsf{E}}$ in $\mathbb{R}^{\mathcal{S}\times\mathcal{A}}/\bm{1}$ for a varying number of expert demonstrations $N^{\textsf{E}}$ and wind strength $\beta$. (c) and (d) show the transferability to $P^{\text{South}}$ and $P^{\text{Shifted}}$ in terms of $\mathsf{SubOpt}_{P^{\text{South}}}(r^{\textsf{E}},\mathsf{RL}_{P^{\text{South}}}(\hat{r}))$ and $\mathsf{SubOpt}_{P^{\text{Shifted}}}(r^{\textsf{E}},\mathsf{RL}_{P^{\text{Shifted}}}(\hat{r}))$, respectively. The circles indicate the median and the shaded areas the 0.2 and 0.8 quantiles over 10 independent realizations of the expert data.